# Improving Quotation Attribution with Fictional Character Embeddings

### v1

@@ 4 Quotation attribution
Non-Explicit Explicit Anaphoric Implicit
SIG $70.0^{+}$
ChatGPT $70.0^{+}$
Unanswerable (%) 20.9 (6.7) 0 24.5 (8.1) 19.8 (8.6)
@@
+ SBERT $70.0^{\star\phantom{\star}}$ (4.5) 98.3 (1.1) $70.8^{\phantom{\star\star}}$ (7.3) $68.1^{\star}$ (5.5)
+ $\text{UAR}_{Reddit}$ $71.1^{\star\star}$ (4.7) 98.7 (1.1) $71.5^{\phantom{\star\star}}$ (7.9) $\mathbf{69.6}^{\star}$ (5.3)
+ $\text{UAR}_{Play}$ $70.9^{\star\phantom{\star}}$ (4.4) 98.7 (1.0) $71.1^{\phantom{\star\star}}$ (7.9) $69.5^{\star}$ (5.7)
+ $\text{UAR}_{Scene}$ $\mathbf{71.2}^{\star\phantom{\star}}$ (4.3) 98.7 (1.0) $\mathbf{71.7}^{\star\star}$ (8.1) $\mathbf{69.6}^{\star}$ (5.4)
@@
In this section, we first review BookNLP’s approach to quotation attribution. BookNLP is a popular Natural Language Processing pipeline dedicated to books. Then, we present our straightforward approach to combine BookNLP’s contextual information with embeddings of characters computed from their explicit quotes and finish by discussing our main results.
4.1 BookNLP
BookNLP’s pipeline includes among others Part-of-speech tagging, Named Entity Recognition, Coreference Resolution and quotation attribution. All components of the pipeline have been trained on LitBank Bamman et al. (2019, 2020); Sims and Bamman (2020), a corpus containing 100 chapters from public domain books. Since its quotation attribution model is not part of a publication, we will review its details in the following paragraphs.
BookNLP carries out quotation attribution in four independents steps: 1) tokenization and named entity recognition, 2) character name clustering, 3) pronominal coreference resolution, and 4) quote-mention linking. The pronominal coreference resolution step outputs mentions along with the predicted character they refer to. Using predicted mentions in the surrounding context of a quote, the quote-mention linking step aims at predicting which mention is the one that refers to the true speaker of the quote. It is achieved by a fine-tuned BERT model Devlin et al. (2019). We give further details of the quote-mention linking step below.
Let $D=(t_{1},\dots t_{n})$ be a tokenized document and $q=(t_{i},\dots,t_{j})$ be a character utterance in this document, starting at token $i$ and ending at token $j$. A contextual segment is computed for each quote, $c_{q}=[\,c_{q}^{left}\;|\,q\,|\;c_{q}^{right}\,]$, where $c_{q}^{left}$ and $c_{q}^{right}$ are left and right contextual information of length $w$ tokens and [|][\,|\,][ | ] denotes concatenation. Additionally, the full quote $q$ is replaced by a special token [QUOTE] and all other quotes occurring in the contextual segment are replaced by the special token [ALTQUOTE], such that the contextual segment mainly reflect narration. Each segment $c_{q}$, containing masked quotes and narrative parts, is fed to BERT, producing a contextual representation $\mathbf{H}_{c_{q}}$. Then an unary compatibility score between the (masked) quote token $q$ and a mention-span $m=(t_{m_{s}},\dots,t_{m_{e}})$ occurring in the segment is calculated as follows:
@@ 4.1 BookNLP
where $\mathbf{H}_{c_{q}}[i]$ is the BERT representation of token $i$ while $|m|$ is the length in tokens of mention $m$ and $\phi$ is a feed-forward neural network with one hidden layer and a tanh activation. Given a set of candidates $Y(q)=(m_{1},\dots,m_{k})$, the model is trained to optimize the log-likelihood of mentions referring to the true speaker of $q$.
The default window size is $w=50$ tokens, and the hidden layer dimension is $100$. During training and inference, only a maximum of $k=10$ mention candidates are considered. In our experiments, we modify these default parameters by increasing the narrative window size to $w=100$ tokens, and using a relu activation and a hidden dimension of $d=512$ in the feed-forward neural network. We also replace BERT by SpanBERT-large Joshi et al. (2020), computing mention representations by concatenating SpanBERT embeddings of the first and last tokens. At inference, we do not truncate the number of candidates and instead use all candidate mentions available in the context window.
@@ 4.2 Integrating Character Embeddings
Let $E(c)=(q_{i},\dots,q_{j})$ be the collection of all explicit quotes uttered by character $c$. We embed this collection of quotes using one of the model described in Section 3.2, yielding a vector representation $v_{c}$ for each character. If a character has not been explicitly quoted, we set its vector representation to zeros. Additionally, we embed each quote $q$ independently with the same model to get a vector representation $u_{q}$. We modify Equation 2 to combine BookNLP’s contextual representation with these character and quote representations:
@@
Equation 3 allows BookNLP to rely on extra-contextual information by injecting global character information and compatibility between quote $q$ and candidate character $c$ derived from CV models. In this work, our primary goal is to prove the usefulness of character embeddings. We leave as future work finding the best strategy to integrate them in systems such as BookNLP.
@@ 4.3 Experimental Setup
Following Su et al. (2023), we use the first version of PDNC, which contains 22 annotated novels. We focus on major and intermediate characters who are characters that uttered at least 10 quotes in the novel. The later allows to avoid the long tail of minor characters and to compare our results to the current state-of-the-art, SIG and ChatGPT.
@@ 4.4 Results
Our main results are displayed in Table 3. Our implementation of BookNLP+ reaches 68.9% accuracy on non-explicit quotes, compared to the 53% reported in Vishnubhotla et al. (2023). Combining BookNLP+ with SBERT character embeddings matches the accuracy of SIG and ChatGPT on non-explicit quotes. The gains are larger for all variants of UAR, reaching 71.2% accuracy on non-explicit quotes with our proposed $\text{UAR}_{Scene}$. As expected, we found that these gains are largely due to the improved attribution performance of implicit quotes, increasing the accuracy by more than 3 points with all UAR models.
@@
Table 4: Variation of accuracy when using predicted speaker of explicit quotes. Statistical significance from paired t-test is indicated with ${}^{\star}=5\%$, ${}^{\star\star}=10\%$
Interestingly, although our proposed variants of UAR tailored for literary texts are better in the CV task, we only see a small – though statistically significant – increase between $\text{UAR}_{Scene}$ and its Reddit counterpart on non-explicit attribution accuracy. The reasons could be twofold. First, BookNLP’s system only focuses on finding the right narrative information that helps understanding the speaker of a quote. However, our embeddings mostly reveal stylistic cues of characters, which are not encoded by BookNLP. Therefore, our implementation provide external character information that might not align perfectly with BookNLP’s contextual decisions, limiting their combination potential. Second, we add character information based on noisy coreference chains, which might hinder the model’s peformance. We believe further research should be conducted in finding better combinations between character embeddings and BookNLP.
The high standard deviations across novels reveals that the task is harder for some novels than others. Looking at the percentage of unanswerable quotes (quotes that do not have a candidate mention referring to their speaker in context) provides a reason for such high variations: it ranges from 0 to 56% per novel. Figure 2 shows the attribution accuracy on non-explicit quotes when only considering novels that have a low percentage of unanswerable quotes. Only 4 novels have a percentage lower than 5%, but 17 out of the 22 have a percentage lower than 20%. As expected, we see higher performance for all models, and the gain from UAR character embeddings is more clear.
@@
[…]
Finally, we present quotation attribution results on 6 annotated novels from the second release of PDNC in Table 5. We use the model trained on the first cross-validation split to predict attributions on each of these novels. The impact of using literary character embeddings is more clear: both of our proposed variants of UAR show increased attribution accuracy against their Reddit counterpart.
5 Conclusion
In this work, we evaluated fictional character embeddings constructed with the explicit quotes they utter, showing that they improve quotation attribution accuracy of non-explicit quotes on PDNC when used in combination with BookNLP’s contextual representations. We created DramaCV, a corpus of English drama plays focused on Character Verification and trained two variants of UAR on this dataset. Our proposed models show better performance when verifying fictional characters on both DramaCV and PDNC. Character embeddings derived from these models are further used for quotation attribution on PDNC, reaching state-of-the-art accuracy when combined with BookNLP+. The impact of using literary character embeddings against embeddings derived from a model trained on Reddit is small, though statistically significant, suggesting superiority of in-domain character representations for quotation attribution.
We found that BookNLP’s attribution accuracy is not distributed uniformly across novels, mainly caused by noisy coreference resolution resulting in many quotes not having candidate mentions that refer to their true speaker. We believe that including our character embeddings in the coreference resolution model as in Li et al. (2023) might improve its performance and thus increase overall attribution accuracy. Another potential solution could include unmasking the closest quotes in the vicinity of a target quote that are likely to contain addressee information. We leave better uses of fictional character embeddings with BookNLP for future exploration.
@@ 6 Limitations
In this work, we focused on integrating fictional character embeddings in encoder-only models. We have not tested with other types of models such as encoder-decoder and decoder only, but we believe that integrating our embeddings in those framework is possible and is likely to improve their performance on the quotation attribution task. Our first set of experiments corroborate this assumption, on the usefulness of our proposed character embeddings in distinguishing fictional characters.
@@
We also proposed models that can be used to create character representations. These models are trained on a corpus of drama plays, which spans almost 400 years of works. It is unclear how the variation of writing style that have occurred throughout these centuries affects the information contained in these embeddings, and how they can generalize to new literary styles. This remark also extends to our results on quotation attribution. We presented results on a corpus containing novels from the 19th to early 20th century ; which is closer in terms of writing style to DramaCV than newer novels published in the 21th century. It is thus unclear how our results generalize to these newer novels, but we strongly believe that accurate character embeddings improve quotation attribution regardless of the writing style of the author.
@@
We only focused on representations encoding stylistic cues and topical preferences and showed that they improve quotation attribution. However, this type of character embeddings might be incomplete and other aspects of characters can be used in addition to our proposed embeddings (personas, gender, objective, or social status). This might allow to build accurate representations of characters, that are likely to improve even more quotation attribution accuracy.
Appendix A Performance on all Test Plays.
@@
Appendix B Computing Information

### v2

@@ 4 Quotation attribution
Motivated by the ability of UAR to encode character utterances in a representation space that contains useful, though insufficient, information for quotation attribution, we evaluate in the next section the impact of combining these global character embeddings with BookNLP’s contextual information. BookNLP is a Natural Language Processing pipeline dedicated to English books, popular among practitioners of Digital Humanities.
4.1 BookNLP
BookNLP’s pipeline includes, among others, Part-of-speech tagging, Named Entity Recognition, Coreference Resolution and quotation attribution. All components of the pipeline have been trained on LitBank Bamman et al. (2019, 2020); Sims and Bamman (2020), a corpus containing 100 chapters from public domain books. Since its quotation attribution model is not part of a publication, we will review its details in the following paragraphs.
BookNLP carries out quotation attribution in four independent steps: 1) tokenization and named entity recognition, 2) character name clustering, 3) pronominal coreference resolution, and 4) quote-mention linking. The first three steps are dedicated to extract named, nominal and pronominal mentions along with the predicted character they refer to. Using extracted mentions in the surrounding context of a quote, the final step, quote-to-mention linking, aims at predicting which mention is the one that refers to the true speaker of the quote. In other words, it finds the information given by the narrator about who is speaking among the narrative parts in the vicinity of quotes. It is achieved by a fine-tuned BERT model Devlin et al. (2019). We give further details of the quote-mention linking step below.
Overall Non-Explicit Explicit Anaphoric Implicit
SIG $72^{+}$ $70.0^{+}$
ChatGPT $71^{+}$ $70.0^{+}$
Unanswerable (%) 14.1 (4.4) 20.9 (6.7) 0 24.5 (8.1) 19.8 (8.6)
@@ 4.1 BookNLP
+ SBERT $79.2^{\star\star}$ (3.8) $70.0^{\star\phantom{\star}}$ (4.5) 98.3 (1.1) $70.8^{\phantom{\star\star}}$ (7.3) $68.1^{\star}$ (5.5)
+ $\text{UAR}_{Reddit}$ $80.1^{\star\star}$ (3.9) $71.1^{\star\star}$ (4.7) 98.7 (1.1) $71.5^{\phantom{\star\star}}$ (7.9) $\mathbf{69.6}^{\star}$ (5.3)
+ $\text{UAR}_{Play}$ $80.0^{\star\phantom{\star}}$ (3.8) $70.9^{\star\phantom{\star}}$ (4.4) 98.7 (1.0) $71.1^{\phantom{\star\star}}$ (7.9) $69.5^{\star}$ (5.7)
+ $\text{UAR}_{Scene}$ $\mathbf{80.2}^{\star\phantom{\star}}$ (3.8) $\mathbf{71.2}^{\star\phantom{\star}}$ (4.3) 98.7 (1.0) $\mathbf{71.7}^{\star\star}$ (8.1) $\mathbf{69.6}^{\star}$ (5.4)
@@
Let $D=(t_{1},\dots t_{n})$ be a tokenized document and $q=(t_{i},\dots,t_{j})$ be a character utterance in this document, starting at token $i$ and ending at token $j$. A contextual segment is computed for each quote, $c_{q}=[\,c_{q}^{left}\;|\,q\,|\;c_{q}^{right}\,]$, where $c_{q}^{left}$ and $c_{q}^{right}$ are left and right contextual information of length $w$ tokens and [|][\,|\,][ | ] denotes concatenation. Additionally, the full quote $q$ is replaced by a special token [QUOTE] and all other quotes occurring in the contextual segment are replaced by a special token [ALTQUOTE], such that the contextual segment mainly reflect narrative parts. Each segment $c_{q}$, containing masked quotes and narrative parts, is fed to BERT, producing a contextual representation $\mathbf{H}_{c_{q}}$. Then an unary compatibility score between the (masked) quote token $q$ and a mention-span $m=(t_{m_{s}},\dots,t_{m_{e}})$ occurring in the segment is calculated as follows:
@@
where $\mathbf{H}_{c_{q}}[i]$ is the BERT representation of token $i$ while $|m|$ is the token length of mention $m$ and $\phi$ is a feed-forward neural network with one hidden layer and a tanh activation. Given a set of candidates $Y(q)=(m_{1},\dots,m_{k})$, the model is trained to optimize the log-likelihood of mentions referring to the true speaker of $q$. Figure 2 summarizes this process.
The default window size is $w=50$ tokens, and the hidden layer dimension is $100$. During training and inference, only a maximum of $k=10$ mention candidates are considered.
In our experiments, we modify these default parameters by increasing the narrative window size to $w=100$ tokens, and using a relu activation and a hidden dimension of $d=512$ in the feed-forward neural network. We also replace BERT by SpanBERT-large Joshi et al. (2020), computing mention representations by concatenating SpanBERT embeddings of the first and last tokens. At inference, we do not truncate the number of candidates and instead use all candidate mentions available in the context window.
@@ 4.2 Integrating Character Embeddings
Let $E(c)=(q_{i},\dots,q_{j})$ be the collection of all explicit quotes uttered by character $c$. We embed this collection of quotes using one of the model described in Section 3.2, yielding a vector representation $v_{c}$ for each character. If a character has not been explicitly quoted, we set its vector representation to zeros. Additionally, we embed each target quote $q$ independently with the same model to get a vector representation $u_{q}$. We modify Equation 2 to combine BookNLP’s contextual representation with these character and quote representations:
@@
Equation 3 allows BookNLP to rely on external information by injecting global character information and compatibility between quote $q$ and candidate character $c$.
@@ 4.3 Experimental Setup
Following Su et al. (2023), we use the first version of PDNC, which contains 22 annotated novels. We focus on major and intermediate characters, who are characters that uttered at least 10 quotes in the novel. This allows to avoid the long tail of minor characters and to compare our results to the current state-of-the-art, SIG and ChatGPT (gpt-3.5-turbo-0613).
@@ 4.4 Results
[…]
Combining BookNLP+ with SBERT character embeddings matches the accuracy of SIG and ChatGPT on non-explicit quotes. The gains are larger for all variants of UAR, reaching 71.2% accuracy on non-explicit quotes with our proposed $\text{UAR}_{Scene}$. As expected, we found that these gains are largely due to the improved attribution performance of implicit quotes, increasing the accuracy by more than 3 points with all UAR models.
Interestingly, although our proposed variants of UAR tailored for literary texts are better at the Authorship Verification of fictional characters, we only see a small – though statistically significant – increase between $\text{UAR}_{Scene}$ and its Reddit counterpart on non-explicit attribution accuracy. The reasons could be twofold. First, BookNLP’s system only focuses on finding the right narrative information that helps identifying the speaker of a quote. However, our embeddings mostly reveal stylistic cues of characters, which are not encoded by BookNLP. Therefore, our implementation provides external character information that might not align perfectly with BookNLP’s contextual decisions, limiting their potential combination. Second, we add character information based on noisy coreference chains, which might hinder the model’s performance. We believe further research should be conducted into finding better combinations between character embeddings and BookNLP.
The high standard deviations across splits reveals that the task is harder for some novels than others. Looking at the percentage of unanswerable quotes (quotes that do not have a candidate mention referring to their speaker in their context window) provides a reason for such high variations: it ranges from 0 to 56% per novel. Figure 3 shows the attribution accuracy of non-explicit quotes when only considering novels that have a low percentage of unanswerable quotes. Only 4 novels have a percentage lower than 5%, but 17 out of the 22 have a percentage lower than 20%. As expected, we see higher performance for all models, and the gain from UAR character embeddings is more clear.
@@
[…]
Finally, we present quotation attribution results for 6 annotated novels from the second release of PDNC in Table 5. We use the model trained on the first cross-validation split to predict attributions in each of these novels. The impact of using literary character embeddings is more clear: both of our proposed variants of UAR show increased attribution accuracy against their Reddit counterpart. Besides, the increase in accuracy over non-explicit quotes is again mostly attributed to an increase in implicit speaker identification.
5 Conclusion
In this work, we evaluated fictional character embeddings constructed with the explicit quotes they utter, showing that they improve quotation attribution accuracy of non-explicit quotes on PDNC when used in combination with BookNLP’s contextual representations. We created DramaCV, a corpus of English drama plays focused on the Authorship Verification of fictional characters and trained two variants of UAR on this dataset. Our proposed models show better performance when distinguishing fictional characters in both DramaCV and PDNC. Character embeddings derived from these models are further used for quotation attribution on PDNC, reaching state-of-the-art accuracy when combined with BookNLP+. The impact of using literary character embeddings against embeddings derived from a model trained on Reddit is small, though statistically significant, suggesting superiority of in-domain character representations for quotation attribution.
@@ 5 Conclusion
We found that BookNLP’s attribution accuracy is not distributed uniformly across novels, mainly caused by noisy coreference resolution, resulting in many quotes not having contextual candidate mentions that refer to their true speaker. We believe that including our character embeddings in the coreference resolution model as in Li et al. (2023) might also improve its performance and thus increase overall attribution accuracy. Another potential solution could include unmasking the closest quotes in the vicinity of a target quote that are likely to contain addressee information. We leave better combination of fictional character embeddings with BookNLP for future exploration.
@@ 6 Limitations
In this work, we focused on integrating fictional character embeddings in encoder-only models. We have not tested with other types of models such as encoder-decoder and decoder only, but we believe that integrating our embeddings in those frameworks is possible and is likely to improve their performance on the quotation attribution task. Our first set of experiments corroborate this assumption, on the usefulness of our proposed character embeddings in distinguishing fictional characters.
@@
We also proposed models that can be used to create character representations. These models are trained on a corpus of drama plays, which spans almost 400 years of work. It is unclear how the variation of writing styles that have occurred throughout these centuries affects the information contained in these embeddings, and how they can generalize to new literary styles. This remark also extends to our results on quotation attribution. We presented results on a corpus containing novels from the 19th to early 20th century ; which is closer in terms of writing style to DramaCV than newer novels published in the 21th century. It is thus unclear how our results generalize to these newer novels, but we strongly believe that accurate character embeddings improve quotation attribution regardless of the writing style of the author.
@@
We only focused on representations encoding stylistic cues and topical preferences and showed that they improve quotation attribution. However, this type of character embedding might be incomplete and other aspects of characters can be used in addition to our proposed embeddings (persona, gender, objective, or social status). This might allow to build accurate representations of characters that are likely to improve even more quotation attribution accuracy.
7 Acknowledgements
This work was partly performed using HPC resources from GENCI-IDRIS and Grid500.
Appendix A Details on the Construction and Evaluation of DramaCV
Each play collected with GutenTag is encoded in a ‘XML’ file, containing various tags such as the <speaker> tag that indicates the speaker of a line. We parse each XML file, removing plays written by non English-speaking authors, and attribute each line to its speaker using the <speaker> tag.
For the Scene split, we automatically check if a play contains a <scene> or <act> tag. We segment plays by default with the <scene> tag, and use the <act> tag if it does not contain a scene tag. We removed plays that did not contain any of these tags. For the Play split, we kept all plays written by English-speaking authors.
In the Authorship Verification terminology, queries and targets are collections of documents written by either the same or different authors. Given a query written by author A and multiple targets written by either the same or different authors, the objective of AV is to recognize all targets written by author A. Details of queries and target construction and our evaluation setup are provided below:
Queries/Targets.
Given characters $(c_{1},\dots,c_{n})$ occurring in a single segment (a scene/act for the Scene split, or the entire play for the Play split), we randomly sample half of their utterances to construct queries $(q_{c_{1}},\dots,q_{c_{n}})$ and use the remaining half to construct targets $(t_{c_{1}},\dots,t_{c_{n}})$. For the train split, we instead randomly sample 8 utterances to build queries and 8 different utterances to build targets.
DramaCV Evaluation.
Using any model $\mathcal{M}$ that is able to encode collections of documents in a feature vector, we encode each query and target independently. Then, we compare the cosine similarity of query vectors with targets: $similarity(\mathcal{M}(q),\mathcal{M}(t))$ $\forall q\in q_{c_{1}},\dots,q_{c_{n}}$ and $\forall t\in t_{c_{1}},\dots,t_{c_{n}}$. For a query $q_{c}$ uttered by character $c$, we calculate an AUC score over the resulting cosine similarities with targets, where a label $1$ is assigned to $t_{c}$ and 00 to other targets. This AUC represents the ability of $\mathcal{M}$ to assign an higher cosine similarities between queries and targets written by the same author. We average this AUC score over each query in a single segment to obtain an AUC score for this particular segment.
PDNC Evaluation.
We use the evaluation setup of Michel et al. (2024) to evaluate our models on the characters of PDNC. This setup proposes two types of evaluation: Character-Character (CC) and Character-Quotes (CQ).
The CC evaluation is similar to the one described above, but the queries are constructed using only the explicit quotes of a character in a single chapter, and targets are constructed using every other utterance in the other chapters. The AUC computed scores how a model is able to assign a higher cosine-similarity between a collection of a few explicit quotes of character, and its other utterances in the novel.
In the CQ setup, queries are similar to CC, but a target is defined as a single utterance in the held-out chapters, without access to contextual information. Therefore, the AUC computed is scoring how similar the collection of a few explicit quotes of character is to its other individual quotes. This task is inherently harder than the CC setup, as an individual quote might be very different from other quotes (e.g “Hey, how are you Lizzy?”).
Appendix B Performance on all Test Plays
@@
Appendix C Computing Information